A manufacturer produces three models of DVD Players in a year. Four times as many of model A are produced as model C, and 7000 more of model B than C. If the total production for the year is 97 000 units, how many of each are produced?
Answer by Cromlix(4381)

Hi there,
Make model C = 'x'
Model A = 4x
Model B = x + 7000
Total production = 97000 units
x + 4x + (x + 7000) = 97000
x + 4x + x + 7000 = 97000
Collect like terms
6x = 90000
x = 15000.
Model C = 15,000 units
Model A = 60,000 units
Model B = 22,000 units
Hope this helps :-)